OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item. : C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (anonymous):

i dont get this at all

OpenStudy (nurali):

C = 0.765x + 0.06(0.765x) X = 100 C = 0.765(100) + 0.06(0.765*100) C = 76.50 + 0.06(76.50) C = 76.50 + 4.59 C = 81.09 (sale price of coat)

OpenStudy (imstuck):

The original price of the item is denoted as "x" in your problem, and 100 is the original cost of the coat Chris is buying. So fill 100 in for every x in your problem and you will get the cost of the coat to Chris.

OpenStudy (imstuck):

Like what Nurali did for you. It's quite a simple concept when you know your values! They gave you 100, so use it where you see the x.

OpenStudy (anonymous):

okay thx so much

OpenStudy (nurali):

My Pleasure.

OpenStudy (tkhunny):

You have to remember that there are two ways to account for multiple discounts. The simplest way is simply to add them together. 15% + 10% = 25% However, this is RARELY used in practice. The much more common method is the sequential method. Start with $100.00 Sale Discount of 15% -- $100.00 * 0.15 = $15.00 Sale Price: $100.00 - $15.00 = $85.00 Employee Discount of 10% -- $85.00 * 0.10 = $8.50 Sale Price: $85.00 - $8.50 = $76.50 <== This is where we get the mysterious "0.765" Tax Rate of 6% -- $76.50 * 0.06 = $4.59 Final Price: $76.50 + $4.59 = $81.09 <== This is where we get the mysterious "0.765" Of course, you could have done this in one giant swoop. $100.00 * (1 - 0.15) * (1 - 0.10) * (1 + 0.06) = $100.00 * 0.85 * 0.90 * 1.06 = 81.09

OpenStudy (nylicia_123):

so what is really the answer